OpenStudy (anonymous):

Write the equation in standard form. - (1/2x + 2y)= 2/3 I've gotten to -1/2x - 2y = 2/3 and I tried multiplying the equation by 3 to get rid of the fractions and it's not working out for me :/

Directrix (directrix):

Is this --> (1/2x supposed to be (1/2) x or 1/(2x) ? I don't know what is in the numerator and what is in the denominator. It makes a major difference in the solution.

OpenStudy (anonymous):

@Directrix It's (1/2)x

Directrix (directrix):

- (1/2) x + 2y= 2/3 Or is it: - [ (-1/2) x + 2y] = 2/3 ----------- Does the negative go just with the (1/2) x or does it also go with the 2y? Once we know that, we can crank out the solution. @dellzasaur

OpenStudy (anonymous):

It's - [ (1/2) x + 2y] = 2/3 (no negative sign on (1/2)x in the original problem) the neg is distributed to both the 1/2 and the 2y I'm pretty sure @Directrix

Directrix (directrix):

First, we need the format for the standard form of a linear equation: Standard Form of a Linear Equation: The standard form of a linear equation is Ax + By = C, where A, B, and C are integers, A> or =0, and A and B are not both zero. SAMPLE Problem: Write y + 5 = (-5/4)( x - 2) in standard form. y + 5 = (-5/4)(x - 2) 4(y + 5) = 4(-5/4)(x - 2) 4y + 20 = -5(x - 2) 4y + 20 = -5x + 10 4y = -5x - 10 5x + 4y = -10 The standard form of the equation is 5x + 4y = -10

Directrix (directrix):

- [ (1/2) x + 2y] = 2/3 -------------------- - (1/2) x - 2y = 2/3 --> distributed the negative Multiply through by (-2) 1x + 4y = -4/3 --> pretty sure answer.

OpenStudy (anonymous):

@Directrix Omggg thank you so much! Really helps a lot. Just for the last answer you gave me, 1x + 4y = -4/3, I'm not allowed to have fractions included in my standard form (my teacher is complicated) haha, that's what I got so confused about.

Directrix (directrix):

I'm not allowed to have fractions included in my standard form ----- 1x + 4y = -4/3 --> pretty sure answer. Let's multiply by 3 for your teacher. 3x + 12y = -4 --> standard form--> no fractions Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers. ----- I hope that negative sign business is correctly stated in the original problem.

OpenStudy (anonymous):

@Directrix Awesome!! Thank you so much. Makes a lot of sense :) I'm pretty sure the negative sign is in it's right place! I really appreciate it! Thanks again!

Directrix (directrix):

Alrighty, then.